OpenStudy (anonymous):

The half-life of Radium-226 is 1590 years. If a sample contains 500 mg, how many mg will remain after 4000 years?

OpenStudy (nurali):

P = P₀*e^(kt). Generally, for half-lives, I like using: P = P₀*2^(kt). Since the half life of Radium-226 is 1590 years, we see that P = (1/2)P₀ at t = 1590. This produces: (1/2)P₀ = P₀*2^(1590k) ==> 1/2 = 2^(1590k) ==> 2^(-1) = 2^(1590k) ==> k = -1/1590. So the require equation is: P = P₀*2^(-t/1590). Therefore, the amount remaining after 4000 years with a 500mg sample is: P = 500 * 2^(-4000/1590) ≈ 87.4mg.